# Word problems and embedding-obstructions in cellular automata groups on groups

### v1

@@ 8 The lamplighter group
Let $C$ be a finite abelian group. The lamplighter group is $L=C\wr\mathbb{Z}=\bigoplus_{\mathbb{Z}}C\rtimes\mathbb{Z}$, where in the latter representation, $\mathbb{Z}$ acts by shifting the coordinates in $\bigoplus_{\mathbb{Z}}C$.
@@
Let $C$ be a finite abelian group, let $A$ be a nontrivial finite alphabet, $n\geq 2$, and $L=C\wr\mathbb{Z}$. Then $\mathrm{Aut}(A^{G})$ has co-NEXPTIME-complete word problems.
@@
Now as the alphabet we take $A=\mathbb{Z}_{5}\times\mathbb{Z}_{6}\times\mathbb{Z}_{7}\times W\times\{0,1,2\}^{3}\times B\times D$, where $B=C^{2}$ and $C$ is the set of colors of Wang tiles, and $D$ is an auxiliary finite set.
@@ 9.2 The PSPACE upper bound
It suffices to show that these exist in the pentagon model. Furthermore, it is easy to reduce the problem to finding a polynomial splitting scheme for “one-sided balls” $P_{k}=\{(x,y)\;|\;y\in[0,k],x\leq 2^{y}\}$, i.e. vertices reachable from $(0,0)$ by only moves that increment the second coordinate. Namely, these sets sandwich contain balls of the pentagon model, of roughly the same radius.
For the set $P_{k}$, a polynomial splitting scheme is obtained from geodesics between $(0,0)$ and vertices of the form $(x,2^{k})$. Specifically, identify $(x,2^{k})$ with $x/2^{k}$ and repeatedly cut the interval in half. ∎
@@ A.2 Polynomial space and alternating polynomial time
In modern CPU design this corresponds roughly to the carry-lookahead adder construction (speeding up the school algorithm of linear carry propagation, which in turn is known as a ripply-carry adder in circuit design); Babbage invented a similar anticipating carriage for his analytical engine.

### v2

@@ 7 The lamplighter group
Let $C$ be a finite nontrivial abelian group. The lamplighter group is $L=C\wr\mathbb{Z}=\bigoplus_{\mathbb{Z}}C\rtimes\mathbb{Z}$, where in the latter representation, $\mathbb{Z}$ acts by shifting the coordinates in $\bigoplus_{\mathbb{Z}}C$.
@@
Let $C$ be a finite nontrivial abelian group, let $A$ be a nontrivial finite alphabet, $n\geq 2$, and $L=C\wr\mathbb{Z}$. Then $\mathrm{Aut}(A^{G})$ has co-NEXPTIME-complete word problems.
@@
Now as the alphabet we take $A=\mathbb{Z}_{5}\times\mathbb{Z}_{6}\times\mathbb{Z}_{7}\times W\times\{0,1,2\}^{2}\times B\times D$, where $B=C^{2}$ and $C$ is the set of colors of Wang tiles, and $D$ is an auxiliary finite set.
@@ 8.2 The PSPACE upper bound
It suffices to show that these exist in the pentagon model. Furthermore, it is easy to reduce the problem to finding a polynomial splitting scheme for “one-sided balls” $P_{k}=\{(x,y)\;|\;y\in[0,k],x\in[0,2^{y}-1]\}$, i.e. vertices reachable from $(0,0)$ by only moves that increment the second coordinate. Namely, these sets sandwich contain balls of the pentagon model, of roughly the same radius.
For the set $P_{k}$, a polynomial splitting scheme is obtained from geodesics between $(0,0)$ and vertices of the form $(x,k)$ with $x\in[0,2^{y}-1]$. Specifically, identify $(x,k)$ with $x/2^{k}$ and repeatedly cut the real interval in half. ∎
@@ A.2 Polynomial space and alternating polynomial time
In modern CPU design this corresponds roughly to the carry-lookahead adder construction (speeding up the school algorithm of linear carry propagation, which in turn is known as a ripple-carry adder in circuit design); Babbage invented a similar anticipating carriage for his analytical engine.